OpenStudy (anonymous):

algebra 1anyone? perform the operation indicated. square root of twelve +3 square root of 75 - 2 and the square root of 27.

OpenStudy (anonymous):

\[\sqrt{12} +3\sqrt{75} - 2\sqrt{27}\]

OpenStudy (anonymous):

√4√3+3(√25)√(3)-2(√9)√3= 2√3+15√3-18√3=-√3 i hope this is the awsner

OpenStudy (anonymous):

my book says the answer is something like :\[ax \sqrt{11a}\]

OpenStudy (anonymous):

any thought of how to get it? @cmath1917

OpenStudy (anonymous):

ok take 4 out of √12 25 out of √3√75 and 9 out of 2√27 then the naturals become 2 5 and 3 respectively then simplify 2√3+3*5√3-2*3√3=11√3 i have changed my last incorrect of -√3